HI, could I get some help with the following problem. The question is: Find the ordered pairs and graph the following equation. x + 2y = 4. Thank You
Answer by checkley71(8403)

USING THE GRAPHING EQUATION Y=mX+b WE GET X+2Y=4 OR 2Y=-X+4 OR Y=(-1/2)X+2 THUS WE HAVE A SLOPE OF -1/2 AND A Y INTERCEPT OF 2
THE ORDERED PAIRS ARE FOUND BY SETTING EACH UNKNOWN (X&Y) TO ZERO AND SOLVING FOR THE OTHER UNKNOWN.
0=(-1/2)X+2 OR (-1/2)X=-2 OR -X=-4 OR X=4 & Y=(-1/2)0+2 OR Y=2
THUS THE ORDERED PAIRS ARE (0,2)(4,0)